# Cross-Modal Geometric Hierarchy Fusion: An Implicit-Submap Driven Framework for Resilient 3D Place Recognition

### v1

@@ Abstract
LiDAR-based place recognition serves as a crucial enabler for long-term autonomy in robotics and autonomous driving systems. Yet, prevailing methodologies relying on handcrafted feature extraction face dual challenges: (1) Inconsistent point cloud density, induced by ego-motion dynamics and environmental disturbances during repeated traversals, leads to descriptor instability, and (2) Representation fragility stems from reliance on single-level geometric abstractions that lack discriminative power in structurally complex scenarios. To address these limitations, we propose a novel framework that redefines 3D place recognition through density-agnostic geometric reasoning. Specifically, we introduce an implicit 3D representation based on elastic points, which is immune to the interference of original scene point cloud density and achieves the characteristic of uniform distribution. Subsequently, we derive the occupancy grid and normal vector information of the scene from this implicit representation. Finally, with the aid of these two types of information, we obtain descriptors that fuse geometric information from both bird’s-eye view (capturing macro-level spatial layouts) and 3D segment (encoding micro-scale surface geometries) perspectives. We conducted extensive experiments on numerous datasets (KITTI, KITTI-360, MulRan, NCLT) across diverse environments. The experimental results demonstrate that our method achieves state-of-the-art performance. Moreover, our approach strikes an optimal balance between accuracy, runtime, and memory optimization for historical maps, showcasing excellent Resilient and scalability. Our code will be open-sourced in the future.
@@ I Introduction
PLACE recognition, the capability of identifying previously visited locations in unmanned systems, serves as a cornerstone for achieving long-term autonomy in robotics and self-driving platforms [1]. This capability is crucial for autonomous mobile robots to achieve precise and robust positioning in unknown environments [2]. Place recognition also has a wide range of applications. For instance, it is used in loop closure detection for Simultaneous Localization and Mapping (SLAM) [3, 4, 5], map merging in large-scale scene mapping [6], and autonomous robot navigation [7, 8]. Light Detection and Ranging (LiDAR) can obtain accurate measurement results even under various lighting conditions, such as in darkness or when disturbed by strong light. As a result, LiDAR plays a crucial role in unmanned systems. Meanwhile, with the rapid development of robotic and autonomous driving systems, there is an increasing demand for LiDAR-based place recognition[9, 10, 11, 12].
Despite the rapid development of place recognition methods based on LiDAR point clouds in recent years, these methods still face the following challenges:
@@
    In an urban environment, the LiDAR point clouds collected at the same location often exhibit inconsistencies in density and coverage due to the variability in the trajectories and motion speeds of vehicles or robots. Therefore, the algorithm must possess variability robust rotational and translational invariance, as well as invariance to point cloud density[13, 14].
@@
    Dynamic objects within the scene, such as cars, pedestrians, and cyclists, continuously alter the 3D structure and distribution of the environment. Their presence is unavoidable and challenges the robustness of LiDAR point cloud-based place recognition [15, 16].
@@
Most existing methods for hand-crafted feature-extraction-based 3D place recognition operate in two primary strategies. One approach is to extract features (such as local[17, 18] or global features[19, 20]) directly from the original 3D point cloud. The other approach involves first converting the raw point cloud into a structured representation (for example, using Bird’s Eye View (BEV)[21], polar coordinate projection[22], intensity images[23], or sparse voxel grids[24]) and then extracting features.
However, existing 3D-based hand-crafted feature extraction methods generally suffer from a fundamental limitation: their reliance on relatively simplistic scene representations. These methods typically focus on feature extraction from single-modal scene representations—either directly from raw point cloud data or by converting point clouds into a specific structured representation (such as BEV or polar coordinates) and relying on it. This dependence on single representations inevitably leads to incompleteness of scene information. The reason is that while raw point clouds preserve precise 3D geometric details, they are susceptible to density variations and lack inherent rotational and translational invariance; structured representations (e.g., BEV), although providing compact global perspectives conducive to performing feature extraction, often sacrifice fine-resolution local geometry or vertical-direction information. Consequently, it is difficult to use features extracted from any single representation to comprehensively and robustly characterize complex 3D scenes, thereby constraining the upper limit of performance for subsequent 3D place recognition tasks.
To address these issues, we have deviated from the previous explicit 3D point cloud input and introduced a point-based implicit representation as the input data. We introduce this representation for three reasons:
@@
    The point-based implicit representation is compact, flexible, and continuous. It possesses rotational and translational invariance, as well as spatial density invariance.
@@
    This representation enables more flexible online removal of dynamic objects, leading to more robust place recognition.recognition.
This paper proposes a brand-new place recognition method based on LiDAR point clouds, as illustrated in Fig. 1. This method uses a point-based implicit representation as the input for scene representation and utilizes the Signed Distance Function (SDF) and stability prediction values to filter out online dynamic objects. Subsequently, it integrates normal and BEV information to enhance place recognition capability. Specifically, we construct a normal-based descriptor by leveraging the angular differences between the surface normals of the key 3D segment. Meanwhile, we generate evenly distributed occupancy grid information using the point-based implicit representation and compute BEV descriptors. Finally, we combine these two types of descriptors to form the final descriptor.
@@
    We introduce a point-based implicit representation. Leveraging its consistent properties and high-quality surface normals, we construct a novel point cloud-based place recognition scheme.
@@
    We propose a cross-modal geometric hierarchy fusion method that synthesizes structural normal descriptors with rotation-invariant BEV features, enabling resilient 3D place recognition through complementary geometric representations.
@@
    We propose a training-free, handcrafted feature extraction method. Our implicit representation enables substantial storage savings and real-time operation. Extensive experiments demonstrate that our method outperforms state-of-the-art approaches.
@@ II Related Works
In this section, we will provide a comprehensive review of implicit map representation, as well as 3D point cloud place recognition methods based on handcrafted and deep learning.
@@ II-A Implicit Neural Map Representation
In recent decades, maps based on explicit representation have been widely applied in unmanned systems[25, 26]. Common methods for explicitly representing scenes include point clouds[27], triangular meshes[28], voxel grids[29], occupancy grids[30], and so on. Recently, the effectiveness of implicit neural representation has been demonstrated in the modeling of radiation fields[31] and geometric fields[32, 33]. Such a representation has facilitated many application scenarios, such as 3D reconstruction, scene generation, novel view synthesis, and so on. Based on NeRF (Neural Radiance Field)[31], DeepSDF[32], and occupancy networks[33], a large amount of research work has been carried out. However, all these works use only a single multi-layer perceptron (MLP) to represent the entire scene. Although this approach is effective, it may face poor scalability, long computation time, and high memory consumption when dealing with complex scenes. Therefore, recent approaches have adopted the hybrid representation to address the above-mentioned issues. This approach jointly optimizes the explicitly stored local latent features and the shallow multi-layer perceptron[34, 35, 36].
With remarkable progress achieved in efficient training methods for implicit neural representations, it has become possible to scale them up to larger scenarios, and they have been widely applied in mapping and Simultaneous Localization and Mapping (SLAM). In the mapping and SLAM systems based on RGB-D data streams, researchers have mainly proposed two types of methods. One is based on a single MLP[37, 38], and the other is based on grid-based local latent features and shallow MLPs[39, 40, 41] to model the geometry or radiance field of the scene while tracking the camera’s pose.
Some work has also been done for the cases where LiDAR is used as the input data stream. IR-MCL[42] and LocNDF[43] proposed a method to construct an implicit neural distance map through laser scanning and to localize the robot within this map. SHINE-mapping[44] aims to extend the implicit neural representation to large-scale outdoor LiDAR data. It stores local features in a sparse voxel format based on an octree, enabling it to handle large-scale scene data while maintaining efficient storage. In the application of SLAM, LONER[45] integrates incremental neural mapping with an ICP-based LiDAR odometry front-end. This approach combines the modeling capabilities of neural networks with traditional point cloud registration algorithms to enhance the accuracy and efficiency of map construction and localization. NeRF-LOAM[46] proposes an implicit neural LiDAR odometry and mapping system based on an online optimizable octree-based feature grid. This method utilizes the concept of NeRF and improves the quality of map construction and the accuracy of localization by optimizing the feature grid. Compared with the previous systems, PIN-SLAM[47] has achieved greater improvements in real-time performance and map storage efficiency. It represents the scene using elastic neural points with fast indexing based on a hash table, enabling the system to operate at a speed close to real-time and saving storage overhead. Inspired by the above work, our work employs a point-based implicit representation. From this representation, we obtain the scene normal vector and occupancy information, which are the foundation for subsequent place recognition.
@@ II-B Handcrafted Methods
Early studies mainly described point clouds by extracting their geometric local features, including distance[48], point normal[49], angle[50], density[51], and so on. These methods have achieved success in point cloud registration and shape recognition. With the in-depth development of the research, researchers have started to attempt to encode the entire point cloud into a global feature descriptor. Z-Projection[52] projects the point cloud onto a certain direction (such as the Z-axis), extracts angular and height information, and then calculates the similarity through the Sørensen distance or the Wasserstein distance. Fast Histogram[53] uses a fast histogram to statistically analyze angular and height information and evaluates the similarity through the Wasserstein distance. M2DP[54] projects the point cloud onto multiple two-dimensional planes, calculates the point density on each plane, and forms a signature vector. This method avoids the calculation of point normals, improving the computational efficiency, but some information may be lost. Scan Context[55, 56] divides the horizontal space into multiple annular regions and fan-shaped regions, forming a two-dimensional height matrix. By performing nearest neighbor search through the annular key and the similarity score, good results have been achieved. To further enhance the performance, researchers have started to combine other types of features, such as intensity[57], spatial binary patterns[58], and so on. LiDAR Iris[59] converts point cloud data into ”iris images”, generates binary signatures through the LoG-Gabor filter and threshold operations, and then calculates the similarity through the Hamming distance. STD[18] proposes a stable triangle descriptor based on the invariance of the triangle shape. By extracting key points and encoding them into triangle descriptors, it enables efficient position matching and geometric verification. Most of the handcrafted methods are relatively natural. However, they generally encounter some problems when faced with large rotations and translations.
@@ II-C Learning-based Methods
PointNetVLAD[17] is a deep learning-based method that combines two crucial components: PointNet[60] and NetVLAD[61]. PointNet is a network that extracts features from point clouds, while NetVLAD is an aggregator that aggregates local features into a 512-dimensional global descriptor. This method lays the foundation for subsequent approaches. Based on PointNetVLAD, two methods, namely SOE-Net[62] and PCAN[63], introduce attention mechanisms. The attention mechanisms can dynamically emphasize the more important parts in the point cloud, thereby enhancing the robustness of the descriptors. Locus[64] enhances the descriptors by introducing temporal information, which can improve the robustness against viewpoint changes. HiTPR[65] partitions the point cloud into dense voxels and utilizes a transformer network[66] to enhance the correlation of local neighbors and the dependence on global context. Consequently, this method demonstrates outstanding performance in highly complex and cluttered environments. Kim et al.[67] proposed a classifier based on Convolutional Neural Network (CNN) for long-term place recognition. This approach can handle significant scene changes over a long time. Semantic Scan Context[68] addresses the translation problem by introducing semantic information. LPD-Net[69] and vLPD-Net[70] utilize Graph Neural Network (GNN) for feature aggregation, thereby effectively capturing the geometric and shape characteristics of the point cloud. DAH-Net[71] proposes a density-driven adaptive hybrid network based on the density changes of point clouds. Through a dynamic local feature aggregation and a contrastive enhanced linear attention module, it attempts to address the challenges posed by the density variations of point clouds in large-scale scenes. These aforementioned learning-based methods often suffer from relatively long training times and require substantial computational resources. Moreover, they exhibit poor performance when faced with scenarios that are not part of their training samples.
@@ III METHOD
Figure 2: Here is the pipeline of our 3D Place Recognition. A specific number of LiDAR frames are fed into form a sub-map, which is then transformed into an implicit neural point representation. High-quality occupancy grid information, triangular points, and normal vector information are subsequently derived from this representation. BEV and primary 3D segments are then obtained. The log-Gabor filter is employed to generate descriptors for the BEV. At the same time, the angular difference distribution of normal vectors is utilized to obtain geometric information descriptors for 3D segments. Finally, these two types of descriptors are fused to create a descriptor for 3D place recognition.
This section will introduce our proposed method for Lidar-based place recognition. In Section III-A, an overview of our method will be presented. Section III-B will demonstrate the neural implicit representation. Section III-D and Section III-E will describe how we obtain descriptors with normal and BEV image information.
@@ III-A Overview
The main process of our proposed method is illustrated in Fig. 2. To better handle the random noise caused by variations in point cloud density and dynamic objects and obtain more discriminative descriptors, we use elastic implicit 3D points to represent the scene.
Considering that the LiDAR point clouds in different scanning modes need to be processed, the input of our method is several frames of point clouds $\{\mathbf{P_{1}},...,\mathbf{P_{n}}\}$. The value of $n$ is determined by calculating the distance traveled by the robot $\tau_{n}$.
Subsequently, for the input point cloud sequence, we use a point-based implicit representation to encode scene geometry. Starting from the first frame, the system generates samples and labels to optimize the neural implicit map, yielding a sub-map representation. From this implicit sub-map, we extract: the scene’s occupancy grid, sample points (mesh vertices) and surface normals at these points.
Next, we remove the road surface from the mesh vertices and apply Euclidean clustering to the remaining area. For each cluster, we design a feature extractor based on angular distributions between adjacent normals, deriving multiple descriptors. Following clustering of these descriptors, we obtain the geometric descriptor $\mathbf{F^{G}}\!\in\!\mathbb{R}^{a}$. Concurrently, we project the occupancy grid into a BEV, transform it to polar coordinates, and compute local responses via Log-Gabor filtering. Keypoints are selected by maximum response, with their descriptors forming the BEV descriptor $\mathbf{F^{B}}\!\in\!\mathbb{R}^{b}$. Finally, we concatenate both descriptors into the fused representation $\mathbf{F^{fuse}}\!\in\!\mathbb{R}^{a+b}$.

### v3

@@ Abstract
LiDAR-based place recognition is critical for long-term autonomous driving without GPS. Existing handcrafted feature methods face dual limitations. First, descriptor instability occurs due to inconsistent point cloud density from motion and environmental changes during repeated traversals. Second, representation fragility arises from reliance on single-level geometric abstractions in complex scenes. To overcome these, we propose a novel framework for 3D place recognition. We introduce an implicit 3D representation using elastic neural points. This method is inherently resilient to variations in input density and yields uniformly distributed representations. From this, we derive occupancy grids and normal vectors. These enable the construction of fused descriptors that integrate complementary perspectives: macro-level spatial layouts from a bird’s-eye view and micro-scale surface geometries from 3D segments. Extensive evaluations on diverse datasets (KITTI, KITTI-360, MulRan, NCLT) demonstrate state-of-the-art performance. Our approach achieves an optimal balance between accuracy, runtime efficiency, and map memory footprint. It exhibits exceptional resilience and scalability. The code will be released soon.
@@ I Introduction
PLACE recognition is the ability to identify previously visited locations in robotics and autonomous driving systems. It serves as a cornerstone for achieving long-term autonomy in transportation systems [1]. This capability is crucial for autonomous mobile robots to achieve precise and robust positioning in unknown environments [2]. Place recognition also has a wide range of applications. For example, it is used in loop closure detection for Simultaneous Localization and Mapping (SLAM) [3, 4, 5], map merging in large-scale scene mapping [6], and autonomous robot navigation [7, 8]. Light Detection and Ranging (LiDAR) can obtain accurate measurement results even under various lighting conditions, such as in darkness or when disturbed by strong light. As a result, LiDAR plays a crucial role in unmanned systems. Meanwhile, with the rapid development of autonomous and robotic driving systems, there is an increasing demand for LiDAR-based place recognition [9, 10, 11, 12].
In recent years, methods for place recognition based on LiDAR point clouds have developed rapidly. However, these methods still face several key challenges in practical applications:
@@
    In urban environments, LiDAR point clouds from the same location often differ in density and coverage. These inconsistencies are due to variations in the trajectories and speeds of vehicles or robots. The algorithm, therefore, must be robust to such variability. It needs to maintain rotational and translational invariance, as well as invariance to point cloud density [13, 14].
@@
    Dynamic objects in the scene, such as cars, pedestrians, and cyclists, constantly change the 3D structure of the environment. Their presence is unavoidable. This poses a challenge to the robustness of LiDAR point cloud-based place recognition [15, 16].
@@
Most current hand-crafted feature extraction-based 3D place recognition methods follow two main strategies. One strategy extracts features (such as local features[17, 18] or global features[19, 20]) directly from the raw 3D point cloud. The alternative technique first transforms the point cloud into a structured format (for example, Bird’s Eye View (BEV)[21], polar coordinate projections[22], intensity images[23], or sparse voxel grids[24]) before extracting features.
However, existing hand-crafted 3D feature extraction However, existing hand-crafted 3D feature extraction methods encounter a fundamental limitation: their dependence on narrow, single-modal scene representations. These methods typically process either unstructured raw point clouds or a particular structured intermediate representation (e.g., BEV, Polar coordinates) derived from the point clouds. Critically, this reliance on any single representation fundamentally undermines feature robustness and coverage. Raw point clouds, although preserving exact geometric details, exhibit non-uniform spatial distributions (density variations) and lack intrinsic rotational and translational invariance, which hampers consistent feature extraction. In contrast, structured representations (e.g., BEV) enable concise global perspectives suitable for feature extraction, but they also forfeit detailed local geometry and vertical context, and perpetuate the point cloud’s spatial irregularities. These inherited irregularities intensify the challenge of obtaining stable and discriminative features. Thus, features extracted solely from one modality remain insufficient for thorough and reliable 3D scene characterization, constraining downstream 3D place recognition performance.
To address these issues, we have moved away from the previous explicit 3D point cloud input and, instead, now use an elastic neural points-based implicit representation. This representation is introduced for three key reasons that are outlined below.
@@
    The elastic neural points-based implicit representation is compact, flexible, and continuous. It possesses rotational and translational invariance, as well as spatial density invariance.
@@
    This representation enables more flexible online removal of dynamic objects, leading to more robust place recognition.
This paper proposes a novel hierarchical geometric fusion approach for LiDAR-based place recognition, as illustrated in Fig. 1. Our method uses a point-based implicit scene representation, combining Signed Distance Function (SDF) and stability prediction to filter dynamic objects online. We synergistically fuse macro-level Bird’s-Eye View (BEV) descriptors, which capture global structural context, with micro-scale normal-based descriptors that encode distinctive local geometry. Specifically, features from BEV descriptors guide the identification and interpretation of salient local features, while the detailed information from normal-based descriptors refines and strengthens the BEV-based scene understanding. This bidirectional guidance enables each descriptor type to enhance the other’s contribution, resulting in a robust and comprehensive scene descriptor that supports enhanced place recognition.
@@
    We introduce a point-based implicit representation. Its consistent properties and high-quality surface normals support a novel point cloud-based place recognition scheme.
@@
    We propose a cross-modal geometric hierarchy fusion method that combines structural normal descriptors with rotation-invariant BEV features. This enables resilient 3D place recognition using complementary geometric representations.
@@
    We propose a training-free, handcrafted feature extraction method. Our implicit representation saves storage and enables real-time operation. Extensive experiments show our method outperforms state-of-the-art approaches.
@@ II Related Works
In this section, we provide a comprehensive review of implicit map representation, as well as 3D point cloud place recognition methods based on both handcrafted and deep learning approaches.
@@ II-A Implicit Neural Map Representation
Explicitly represented maps (e.g., point clouds[27], meshes[28], voxel/occupancy grids[29, 30]) have been widely used in unmanned systems[25, 26]. Recently, implicit neural representations have demonstrated effectiveness in modelling radiation[31] and geometric fields[32, 33], enabling applications such as 3D reconstruction and novel view synthesis. Methods such as NeRF[31], DeepSDF[32], and occupancy networks[33] typically employ a single MLP to represent entire scenes. While effective, this approach faces efficiency bottlenecks (scalability, computation, and memory) in complex scenes. To address this, recent work has adopted hybrid representations that jointly optimize explicitly stored local latent features with shallow multilayer perceptrons (MLPs) [34, 35, 36].
Advances in efficient training now enable the scaling of implicit neural representations to larger scenes, accelerating their adoption in mapping and SLAM. For RGB-D SLAM, two main approaches exist for simultaneous scene modeling and camera pose tracking: one uses a single MLP to represent the entire scene as a global implicit function [37, 38], while the other uses grid-based local latent features with multiple shallow MLPs, enabling local modelling and improved scalability[39, 40, 41].
In addition to RGB-D approaches, LiDAR-based systems have also advanced, with methods like IR-MCL[42] and LocNDF[43] constructing implicit neural distance maps for robot localization. Building upon these developments, SHINE-mapping [44] extends implicit neural representations to large-scale outdoor LiDAR data using octree-based sparse voxel features for efficient storage. LONER[45] integrates incremental neural mapping with an ICP-based LiDAR odometry front-end for continuous scene updates. NeRF-LOAM[46] implements an online-optimizable octree feature grid, enabling efficient LiDAR odometry and mapping. PIN-SLAM[47] employs elastic neural points with hash-table indexing, significantly improving real-time performance and storage efficiency over previous neural mapping techniques. These LiDAR-based works demonstrate complementary advantages of implicit neural representations, including spatial continuity and rich feature encoding. Inspired by these distinctions, our method employs a point-based implicit neural representation. From this continuous representation, we directly derive scene surface normals and occupancy information, forming the essential basis for subsequent place recognition.
@@ II-B Handcrafted Methods
Early studies primarily described point clouds by extracting their geometric local features, including distance [48], point normal [49], angle [50], and density [51], among others. These methods have achieved success in point cloud registration and shape recognition. As research progressed, researchers began to encode the entire point cloud into a global feature descriptor. For example, Z-Projection [52] projects the point cloud onto a certain direction (such as the Z-axis), extracts angular and height information, and then calculates the similarity through the Sørensen distance or the Wasserstein distance. Similarly, Fast Histogram [53] utilizes a fast histogram to statistically analyze angular and height information, and evaluates similarity using the Wasserstein distance. Another approach, M2DP [54], projects the point cloud onto multiple two-dimensional planes, calculates the point density on each plane, and forms a signature vector. This method avoids calculating point normals, thereby improving computational efficiency; however, some information may be lost. In contrast, Scan Context [55, 56] divides the horizontal space into multiple annular regions and fan-shaped regions, forming a two-dimensional height matrix. By performing nearest neighbor search through the annular key and the similarity score, good results have been achieved. To further enhance performance, researchers have begun to combine other types of features, such as intensity[57] and spatial binary patterns[58], among others. For instance, LiDAR Iris[59] converts point cloud data into ”iris images”, generates binary signatures through the LoG-Gabor filter and threshold operations, and then calculates the similarity through the Hamming distance. In another direction, STD[18] proposes a stable triangle descriptor based on the invariance of the triangle shape. By extracting key points and encoding them into triangle descriptors, it enables efficient position matching and geometric verification. Additionally, BoW3D[82] employs an innovative bag-of-words model to process 3D LiDAR features. Finally, RING++[83] constructs a rotation and translation invariant representation with strict mathematical guarantees by Radon transform, Fourier transform, and cross correlation, and solves the global localization problem on sparse maps.
Most handcrafted feature methods utilize a single modality (e.g., point clouds, voxels, BEV), resulting in information loss as cues from other sources are overlooked. While simple, these methods struggle with major rigid transformations or complex scenes. To address this, our approach uses an implicit neural point representation that combines 3D segment normals and BEV features. By hierarchically fusing these modalities, we build a more robust and comprehensive representation for challenging scenarios.
@@ II-C Learning-based Methods
PointNetVLAD[17] is a deep learning-based method that combines two crucial components: PointNet[60] and NetVLAD[61]. PointNet extracts features from point clouds. NetVLAD aggregates these local features into a 512-dimensional global descriptor. This method lays the foundation for subsequent approaches. Building on PointNetVLAD, two methods—SOE-Net[62] and PCAN[63]—introduce attention mechanisms. These mechanisms dynamically emphasise the more important parts in the point cloud, enhancing the robustness of the descriptors. Locus[64] enhances the descriptors by introducing temporal information. This addition improves robustness against viewpoint changes. HiTPR[65] partitions the point cloud into dense voxels and uses a transformer network[66] to enhance correlation among local neighbours and dependency on global context. As a result, HiTPR demonstrates outstanding performance in highly complex and cluttered environments. Kim et al.[67] proposed a classifier based on a Convolutional Neural Network (CNN) for long-term place recognition. This approach can handle significant scene changes over an extended period. Semantic Scan Context [68] addresses the translation problem by introducing semantic information. LPD-Net[69] and vLPD-Net[70] use Graph Neural Networks (GNN) for feature aggregation. This enables them to capture geometric and shape characteristics of the point cloud effectively. DAH-Net[71] proposes a density-driven adaptive hybrid network using density changes of point clouds. It combines dynamic local feature aggregation with a contrastively enhanced linear attention module. With these, it attempts to address the challenges caused by density variations in point clouds from large-scale scenes. LCDNet [84] introduces a novel, unbalanced, optimal transport-based differentiable module. This module enables joint loop closure detection and 6-DoF pose estimation, allowing robust registration under arbitrary initial rotations and excelling at reverse loops. BEVPlace++[85] introduces a Rotation Equivariant Module (REM), enabling weakly supervised, real-time LiDAR global localisation. It extracts rotation-equivariant features from Bird’s Eye View images, achieving robust 3-DoF pose estimation without precise ground truth. These learning-based methods often require relatively long training times and substantial computational resources. Moreover, they perform poorly when facing scenarios that fall outside their training samples.
@@ III METHOD
Figure 2: This is the pipeline for our 3D Place Recognition. We feed a specific number of LiDAR frames to form a sub-map, which we then transform into an implicit neural point representation. From this, we derive high-quality occupancy grid information, triangular points, and normal vectors. BEV and primary 3D segments are obtained next. We use a log-Gabor filter to generate descriptors for the BEV, while the angular difference distribution of normal vectors yields geometric descriptors for the primary 3D segments. Finally, we combine these two descriptor types to form a single one for 3D place recognition. In the figure, yellow cuboids show macro descriptors from the bird’s-eye view, and blue cuboids show micro descriptors from the main 3D segments. The two descriptors are hierarchically fused to integrate their advantages.
This section will introduce our proposed method for Lidar-based place recognition. In Section III-A, an overview of our method will be presented. Section III-B will demonstrate the elastic neural implicit representation. Section III-C and Section III-D will describe how we obtain hierarchy fusion descriptors with normal and BEV image information.
@@ III-A Overview
The main process of our proposed method is illustrated in Fig. 2. To better handle the random noise caused by variations in the density of point clouds (which refers to fluctuations in how densely the 3D data points, collected by sensors, are distributed in space) and by dynamic objects (objects in the environment that move or change over time), and to obtain more discriminative descriptors (distinctive features used to tell apart different parts of a scene), we use elastic neural implicit 3D points. This representation models the environment as a continuous field using a neural network, allowing for the flexible positioning of points to more effectively describe the scene.
Because LiDAR (a technology that measures distances using laser light to produce detailed 3D maps) point clouds acquired in different scanning modes need to be processed, our method takes as input several frames of point clouds $\mathbf{P_{1}},...,\mathbf{P_{n}}\$$}, where each frame is a collection of 3D data points. The number $n$ depends on the distance the robot has travelled, denoted by $\tau_{n}$, which represents the path length corresponding to $n$ frames.
Next, for the input sequence of point cloud frames, we utilise a point-based implicit representation to encode the scene’s geometry. Starting from the first frame, the system generates samples (selected points) and labels (information or categories assigned to samples) to optimise the neural implicit map, which results in a sub-map (a smaller localised map) representation. We then extract from this implicit sub-map: the scene’s occupancy grid (a division of space into a grid where each cell indicates occupancy), sample points (vertices of a generated mesh approximating surfaces), and surface normals (vectors perpendicular to the surface at each mesh vertex).
Then, we use Euclidean clustering (a technique that groups points based on how close they are to each other in 3D space) to quickly extract the main 3D segments of the scene. These segments guide the projection of the occupancy grid and the creation of the final bird’s-eye-view (BEV) representation—a top-down 2D view of the 3D space. We detect keypoints (distinct and important points in the data) on the BEV map, and calculate their spatial statistics (measures that summarise their distribution). These statistics provide initial information used to guide the extraction of feature descriptors (numeric vectors that describe local geometric properties) in the main 3D segments, thus generating a dense set of micro-scale surface signatures $\mathbf{F^{G}}\in\mathbb{R}^{a}$. Next, we aggregate the BEV-derived features using unsupervised clustering (a method to find groups in unlabeled data), to produce a compact BEV descriptor $\mathbf{F^{B}}\in\mathbb{R}^{b}$ that is highly distinctive. Finally, both descriptors are combined into a single fused feature representation $\mathbf{F^{fuse}}\in\mathbb{R}^{a+b}$.